OpenStudy (anonymous):

USE method of substution to solve the system of linear equatons. 4y-y=8 9x-2y=20 Solution? type order pair could some one show me how to get the answer on this

OpenStudy (mathlegend):

What I would do first is subtract the 4y from the first equation...

OpenStudy (mathlegend):

-y= -4y + 8

OpenStudy (anonymous):

4x-y=8 9x-2y=20

OpenStudy (mathlegend):

Then divide by -1

OpenStudy (anonymous):

use the top equation to solve for y and replace that value for the y in the second equation

OpenStudy (mathlegend):

Now, since we solved for y... we can substitute it into the second equation as @emilie.li said.

OpenStudy (mathlegend):

9x-2(4x-8)=20 9x-8x+16=20 x+16=20 x= 4

OpenStudy (mathlegend):

That problem up there should say y=4x-8

OpenStudy (anonymous):

to be honest I don't know how to do any of this math homework

OpenStudy (mathlegend):

@Ne1977 do you understand how I solved for x?

OpenStudy (mathlegend):

@Ne1977 I need to know do you understand how I solved for 'x'? I will solve for 'y' if you understand.

OpenStudy (anonymous):

yes I got 4, and 8

OpenStudy (mathlegend):

@Ne1977 you are so right! :)

OpenStudy (mathlegend):

Good job!